OpenStudy (anonymous):

Multiply. (b+frac{1}{3})(b+frac{1}{5})=

OpenStudy (anonymous):

\[(b+ \frac{1}{3})(b+\frac{1}{5})\]

OpenStudy (anonymous):

\[b^2 + \frac{b}{5}+\frac{b}{3} + \frac{1}{15}\] \[b^2 + \frac{8b}{15} +\frac{1}{15}\]